# Fact-checking information generated by a large language model can decrease news discernment

### v3

@@ Appendix 3 Regression analyses
In our preregistered research design, we proposed an exploratory analysis employing logistic cross-classified multilevel modeling (MLM) to predict item-level response accuracy. This approach categorizes responses into two distinct groups: those considered desirable (i.e., believing or sharing true news, and not believing or sharing false news) and those deemed undesirable (believing or sharing false news, and not believing or sharing true news). However, we later noticed two problems with this approach that drove us to pursue a different exploratory analysis. First, the MLM experienced issues converging properly, raising doubts about its reliability. Second, we recognized that this methodology does not allow us to separately analyze responses to true and false news, crucial to assessing discernment. Consequently, to align with the analysis in the main text, we opted to employ linear regression with clustered standard errors, focusing on participant responses as the dependent variable. This deviation from our preregistered exploratory design brings our methodology in line with precedents set in the literature50, 70. The dependent variable in all models is the participant’s response indicating belief or willingness to share a specific headline, coded as 1 for “Yes” and 0 for “No.” Age and Education level (as described in Education) are included as covariates in all analyses.
@@ 3.1 Ineffectiveness of LLM fact checks
To examine the robustness of the finding that exposure to LLM fact checks did not significantly affect participants’ discernment, our model incorporates dummy variables for the experimental Condition and headline Veracity, as well as a variable for their interaction51, 50.
Tables S1 and S2 display the results obtained from fitting our data to this model for the belief and share groups, respectively. Of particular relevance to our primary findings, the interaction terms of interest, namely “Condition(Forced):Veracity(True)” and “Condition(Optional):Veracity(True),” are not significant predictors in either model. As shown previously51, 50, the coefficients of such interaction terms directly quantify the average change in discernment driven by each respective experimental treatment. Thus, this analysis provides further support for our result that exposure to LLM fact-checking information did not significantly impact average discernment.
Table S1: Ineffectiveness of LLM Fact Checks Coefficients (Belief Group; $F=971.64$, $R^{2}=0.19$, $P<0.001$)
@@
(Intercept) 0.563 0.050 11.229 $<0.001$**
Condition(Forced) 0.035 0.030 1.180 0.238
Condition(Optional) 0.007 0.030 0.220 0.826
Veracity(True) 0.393 0.027 14.774 $<0.001$**
Age -0.006 0.001 -7.299 $<0.001$**
Education 0.008 0.006 1.358 0.175
@@
Table S2: Ineffectiveness of LLM Fact Checks Coefficients (Share Group; $F=561.35$, $R^{2}=0.11$, $P<0.001$)
@@
(Intercept) 0.835 0.043 19.449 $<0.001$**
Condition(Forced) 0.049 0.029 1.706 0.088 $\cdot$
Condition(Optional) -0.006 0.031 -0.195 0.846
Veracity(True) 0.085 0.018 4.708 $<0.001$**
Age -0.008 0.001 -10.704 $<0.001$**
Education -0.022 0.009 -2.528 0.011 *
Condition(Forced):Veracity(True) -0.004 0.017 -0.262 0.794
@@ 3.2 Accounting for LLM accuracy
Tables S3 and S4 present the results of fitting our data to this model for the belief and share groups, respectively. [^1] Note that a few standard errors cannot be computed leading to ‘NaN’ values in Tables S4, S13, and S14. This occurs only in the “False $\times$ unsure” scenario, likely due to the low number of headlines (two) in that scenario. Some significant interaction terms are observed for specific FC scenarios. This tells us that the Condition $\times$ FC Scenario relationship is significantly different in these scenarios relative to the “reference group” FC Scenario (False $\times$ false) — not shown in the table. However, this is not the appropriate reference group: we wish to specifically test the significance of this interaction within each fact-checking scenario. To this end, we conduct post-hoc comparisons similar to those presented within the main text for each group. Utilizing the fitted models, estimated marginal mean values for the Control and Forced groups are calculated and compared in each headline scenario, adjusting $P$ values with Bonferroni’s method. The results of these post-hoc comparisons for the belief and share groups are shown in Tables S5 and S6, respectively. We observe significant mean differences for fact-checking scenarios in both groups that are consistent with those presented in the main text. However, in the share group, we also observe a significant mean difference when False headlines were correctly identified as False by the LLM, suggesting that the LLM fact-checking information is harmful in this case as well. This last result is not reported in the main text because it is inconsistent with the corresponding analysis based on mean differences calculated from the raw data, where adjusted $P=1$ suggests insignificant effects.
@@
Cond.(Forced):FC Scen.(False $\times$ unsure) 0.070 0.024 2.865 0.004 **
@@
Headline Scenario Control -- Forced Std. Err. df $t$ ratio Adj. $P^{\dagger}$ Sig.
True $\times$ False 0.120 0.020 19508 5.915 $<$ 0.001 ***
True $\times$ Unsure -0.012 0.011 19508 -1.021 1.000
True $\times$ True -0.048 0.023 19508 -2.063 1.000
False $\times$ False -0.029 0.010 19508 -2.995 0.124
False $\times$ Unsure -0.098 0.029 19508 -3.427 0.028 *
@@
$\dagger$ Bonferroni’s method comparing a family of 45 estimates
@@
Headline Scenario Control -- Forced Std. Err. df $t$ ratio Adj. $P^{\dagger}$ Sig.
True $\times$ False -0.015 0.020 21428 -0.756 1.000
True $\times$ Unsure -0.045 0.011 21428 -3.992 0.003 **
True $\times$ True -0.091 0.023 21428 -3.922 0.004 **
False $\times$ False -0.047 0.001 21428 -4.943 $<$ 0.001 ***
False $\times$ Unsure -0.078 0.028 21428 -2.739 0.277
@@
$\dagger$ Bonferroni’s method comparing a family of 45 estimates
@@ 3.3 Opt in versus opt out
To provide support for our analysis related to the optional condition, we now incorporate a dummy variable that captures whether a participant in this condition chose to see LLM fact-checking information (opt in) or not (opt out).
Tables S7 and S8 present the results of fitting our data for the belief and share groups, respectively. To confirm the results presented in the main text, we utilize the models to perform the same comparisons of estimated marginal means. These post-hoc comparisons further support our findings, and are shown for the belief and share groups in Tables S9 and S10, respectively.
Table S7: Opt In versus Opt Out Coefficients (Belief Group; $F=571.07$, $R^{2}=0.21$, $P<0.001$)
@@
(Intercept) 0.637 0.061 10.523 $<0.001$**
Option(opt out) -0.245 0.038 -6.406 $<0.001$**
Veracity(True) 0.259 0.039 6.671 $<0.001$**
Age -0.004 0.001 -3.817 $<0.001$**
Education -0.008 0.009 -0.846 0.398
Option(opt out):Veracity(True) 0.267 0.047 5.679 $<0.001$**
@@
Table S8: Opt In versus Opt Out Coefficients (Share Group; $F=472.78$, $R^{2}=0.18$, $P<0.001$)
@@
(Intercept) 0.806 0.066 12.159 $<0.001$**
Option(opt out) -0.306 0.038 -7.949 $<0.001$**
Veracity(True) 0.058 0.017 3.323 $<0.001$**
Age -0.007 0.001 -5.255 $<0.001$**
Education 0.001 0.014 0.054 0.957
Option(opt out):Veracity(True) 0.018 0.024 0.749 0.454
@@
Veracity Opt in -- Opt out Std. Error df $t$ ratio Adj. $P^{\dagger}$ Sig.
True -0.022 0.013 10434 -1.735 0.4964
False 0.245 0.013 10434 19.371 $<$ 0.001 ***
@@
$\dagger$ Bonferroni’s method comparing a family of 6 estimates
@@
Veracity Opt in -- Opt out Std. Error df $t$ ratio Adj. $P^{\dagger}$ Sig.
True 0.287 0.013 10514 22.525 $<0.001$**
False 0.306 0.013 10514 24.095 $<0.001$**
@@
$\dagger$ Bonferroni’s method comparing a family of 45 estimates
@@ Appendix 4 Interaction analyses
In this section, we explore the potential moderation effects of two factors on our main results: attitude towards AI (ATAI) and headline congruence (see the Covariates section for details). We employ linear regression with robust standard errors clustered on participant and headline for each key finding discussed in the main text. Each analysis covered in Regression analyses is revisited to incorporate these variables and create three-way interactions. Covariates that were included in the earlier analyses (Age and Education level) are included again. The belief and sharing group data are modeled separately.
@@ 4.1 Attitude towards AI
We begin by examining whether LLM fact-checking information remains ineffective amongst individuals with varying levels of ATAI. Therefore, we test the three-way interaction between Condition, Veracity, and ATAI (Condition $\times$ Veracity $\times$ ATAI). Figure S1 presents the relationship between participants’ ATAI and their belief in (panel a) and intent to share (panel b) true versus false headlines across all conditions. The results of our modeling analysis indicate that there is no significant three-way interaction between ATAI and either belief in (Table S11) or intent to share (Table S12) headlines for all conditions.
@@
Condition(Forced):Veracity(True) 0.027 0.085 0.314 0.753
@@
Condition(Optional):Veracity(True) 0.087 0.066 1.321 0.186
@@
Next, we examine whether the effects of fact-checking scenarios stay consistent among people with different ATAI (the three-way interaction Condition $\times$ FC Scenario $\times$ ATAI). Again, we focus on the forced and control conditions and exclude data for the optional participants when fitting each model. Figure S2 illustrates the relationship between belief in headlines and ATAI for the control and forced conditions in each fact-checking scenario. The same relationship is presented with respect to sharing intent in Figure S3. The result of fitting the belief and share group models are found in Tables S13 and S14, respectively. These models are then utilized for post-hoc comparisons similar to those presented within the main text for each group. However, to test for an ATAI interaction, we now compare the slopes of the Control and Forced groups predicted response line, given different values of ATAI. These results, shown in Tables S15 and S16 for the belief and share groups, respectively, validate our results by illustrating that participants did not respond differently depending on ATAI.
@@
Headline Scenario Control -- Forced Std. Err. df $t$ ratio Adj. $P^{\dagger}$ Sig.
False $\times$ false -0.010 0.008 19498 -1.234 1.000
False $\times$ unsure -0.023 0.023 19498 -1.000 1.000
True $\times$ false 0.038 0.016 19498 2.313 0.934
True $\times$ unsure $-1.234\times 10^{-4}$ 0.009 19498 -0.013 1.000
True $\times$ true -0.013 0.019 19498 -0.658 1.000
@@
$\dagger$ Bonferroni’s method comparing a family of 45 estimates
@@
Headline Scenario Control -- Forced Std. Err. df $t$ ratio Adj. $P^{\dagger}$ Sig.
False $\times$ false 0.013 0.008 21418 1.537 1.000
False $\times$ unsure -0.012 0.025 21418 -0.495 1.000
True $\times$ false $2.957\times 10^{-4}$ 0.018 21418 0.017 1.000
True $\times$ unsure 0.014 0.010 21418 1.449 1.000
True $\times$ true -0.015 0.021 21418 -0.746 1.000
@@
$\dagger$ Bonferroni’s method comparing a family of 45 estimates
Next, we examine whether behavior in the optional condition depends on ATAI by introducing a three-way interaction term involving whether a participant chose to view LLM fact checks (opt in vs. opt out), headline veracity, and individual attitude towards AI (Opt-Condition $\times$ Veracity $\times$ ATAI). The results for the belief and share groups are presented in Tables S17 and S18, respectively. We conduct a post-hoc analysis that compares the slopes of the opt-in and opt-out conditions across varying levels of ATAI for the belief (Table S19) and sharing (Table S20) groups, respectively. Results of the post-hoc comparisons can be found in Tables S21 and S22 for the belief and sharing groups, respectively.
@@
Table S17: Opt In versus Opt Out Coefficients (ATAI interaction, Belief Group; $F=319.53$, $R^{2}=0.22$, $P<0.001$)
@@
(Intercept) 0.602 0.122 4.921 $<0.001$**
Option(opt out) -0.197 0.120 -1.643 0.100
Veracity(True) 0.307 0.077 3.982 $<0.001$**
ATAI 0.008 0.023 0.355 0.723
Age -0.004 0.001 -3.900 $<0.001$**
Education -0.007 0.009 -0.718 0.473
Option(opt out):Veracity(True) 0.332 0.109 3.063 0.002 **
Option(opt out):ATAI -0.010 0.026 -0.390 0.697
Veracity(True):ATAI -0.010 0.014 -0.714 0.475
Option(opt out):Veracity(True):ATAI -0.017 0.022 -0.765 0.444
@@
(Intercept) 0.613 0.143 4.295 $<0.001$**
Option(opt out) -0.152 0.142 -1.076 0.282
Veracity(True) 0.048 0.058 0.832 0.406
ATAI 0.041 0.027 1.518 0.129
Age -0.007 0.001 -5.055 $<0.001$**
Education -$3.125\times 10^{-4}$ 0.014 -0.022 0.982
Option(opt out):Veracity(True) 0.052 0.087 0.591 0.555
Option(opt out):ATAI -0.032 0.030 -1.055 0.291
Veracity(True):ATAI 0.003 0.013 0.200 0.842
Option(opt out):Veracity(True):ATAI -0.008 0.019 -0.430 0.667
@@
Option Veracity $b$ Std. Err. df $t$-ratio $P$ Sig.
Opt in False 0.008 0.006 10430 1.250 0.211
Opt out False -0.002 0.007 10430 -0.301 0.763
Opt in True -0.002 0.006 10430 -0.358 0.720
Opt out True -0.029 0.007 10430 -3.892 $<$ 0.001 ***
@@
Condition Veracity $b$ Std. Err. df $t$-ratio $P$ Sig.
Opt in False 0.041 0.007 10510 5.881 $<$ 0.001 ***
Opt out False 0.009 0.008 10510 1.125 0.260
Opt in True 0.043 0.007 10510 6.551 $<$ 0.001 ***
Opt out True 0.003 0.008 10510 0.390 0.700

### v4

@@ Appendix 3 Regression analyses
In our preregistered research design, we proposed an exploratory analysis employing logistic cross-classified multilevel modeling (MLM) to predict item-level response accuracy. This approach categorizes responses into two distinct groups: those considered desirable (i.e., believing or sharing true news, and not believing or sharing false news) and those deemed undesirable (believing or sharing false news, and not believing or sharing true news). However, we later noticed two problems with this approach that drove us to pursue a different exploratory analysis. First, the MLM experienced issues converging properly, raising doubts about its reliability. Second, we recognized that this methodology does not allow us to separately analyze responses to true and false news, crucial to assessing discernment. Consequently, to align with the analysis in the main text, we opted to employ linear regression with clustered standard errors, focusing on participant responses as the dependent variable. This deviation from our preregistered exploratory design brings our methodology in line with precedents set in the literature(51, 81). The dependent variable in all models is the participant’s response indicating belief or willingness to share a specific headline, coded as 1 for “Yes” and 0 for “No.” Age and Education level (as described in Education) are included as covariates in all analyses. Participants’ age is scaled by a factor of 10 to facilitate the interpretation of the coefficients, allowing for a more straightforward understanding of the effects associated with each decade of age rather than each individual year. Finally, for the sake of brevity, we sometimes use “Optional” and “Forced” interchangeably with “LLM-optional” and “LLM-forced” to describe these experimental conditions.
@@ 3.1 Ineffectiveness of LLM fact checks
To examine the robustness of our findings related to the effects of different treatments on participants’ average discernment, our model incorporates dummy variables for the experimental conditions and headline veracity, as well as a term accounting for their interactions(52, 51).
Tables S3 and S4 display the results obtained from fitting our data to this model for the belief and share groups, respectively. Of particular relevance to our primary findings, the interaction terms of interest, namely “Condition(Forced):Veracity(True)” and “Condition(Optional):Veracity(True),” are not significant predictors in either model. On the other hand, the “Condition(HumanFC):Veracity(True)” interaction term is significant in both models. As shown previously(52, 51), the coefficients of such interaction terms directly quantify the average change in discernment driven by each respective experimental treatment. Therefore, this analysis reinforces our finding that exposure to LLM fact-checking information did not significantly affect average discernment, whereas human fact checks led to an increase in average discernment.
Table S3: Ineffectiveness of LLM Fact Checks Coefficients (Belief Group; $F=1454.23$, $R^{2}=0.24$, $P<0.001$)
@@
(Intercept) 0.545 0.048 11.266 $<0.001$**
Condition(Forced) 0.035 0.030 1.160 0.246
Condition(Optional) 0.005 0.030 0.183 0.855
Condition(HumanFC) -0.011 0.029 -0.379 0.705
Veracity(True) 0.393 0.027 14.776 $<0.001$**
Age -0.006 0.001 -7.250 $<0.001$**
Education 0.009 0.005 1.909 0.056 $\cdot$
@@
Table S4: Ineffectiveness of LLM Fact Checks Coefficients (Share Group; $F=599.84$, $R^{2}=0.11$, $P<0.001$)
@@
(Intercept) 0.826 0.041 20.392 $<0.001$**
Condition(Forced) 0.049 0.029 1.698 0.090 $\cdot$
Condition(Optional) -0.008 0.031 -0.262 0.794
Condition(HumanFC) 0.020 0.029 0.682 0.495
Veracity(True) 0.085 0.018 4.710 $<0.001$**
Age -0.008 0.001 -13.980 $<0.001$**
Education -0.016 0.007 -2.389 0.017 *
Condition(Forced):Veracity(True) -0.004 0.017 -0.262 0.793
@@ 3.2 Accounting for LLM accuracy
Tables S5 and S6 present the results of fitting our data to this model for the belief and share groups, respectively. [^1] Note that a few standard errors cannot be computed leading to ‘NaN’ values. This occurs only for some terms related to the “False $\times$ unsure” scenario, likely due to the low number of headlines (two) in that scenario. Some significant interaction terms are observed for specific FC scenarios. This tells us that the Condition $\times$ FC Scenario relationship is significantly different in these scenarios relative to the “reference group” FC Scenario (False $\times$ false)—not shown in the table. However, this is not the appropriate reference group: we wish to specifically test the significance of this interaction within each fact-checking scenario. To this end, we conduct post-hoc comparisons similar to those presented within the main text for each group. Utilizing the fitted models, estimated marginal mean values for the Control and Forced groups are calculated and compared in each headline scenario, adjusting $P$ values with Bonferroni’s method. The results of these post-hoc comparisons for the belief and share groups are shown in Tables S7 and S8, respectively. We observe significant mean differences for fact-checking scenarios in both groups that are consistent with those presented in the main text. However, we also observe significant mean differences suggesting that the LLM fact-checking information is harmful in additional fact checking scenarios within both the belief (False $\times$ false) and sharing (False $\times$ false, False $\times$ unsure, and True $\times$ unsure) groups. To remain conservative in our analyses, we do not report these results in the main text as they are inconsistent with the corresponding analysis based on mean differences calculated from the raw data.
@@
Cond.(Forced):FC Scen.(False $\times$ unsure) 0.070 0.024 2.864 0.004 **
@@
Headline Scenario Forced -- Control Std. Err. df $t$ ratio Adj. $P^{\dagger}$ Sig.
True $\times$ False -0.120 0.020 19508 -5.915 $<$ 0.001 ***
True $\times$ Unsure 0.012 0.011 19508 1.021 1.000
True $\times$ True 0.048 0.023 19508 2.063 0.196
False $\times$ False 0.029 0.010 19508 2.995 0.014 *
False $\times$ Unsure 0.098 0.029 19508 3.427 0.003 **
@@
$\dagger$ Bonferroni’s method comparing a family of 5 estimates
@@
Headline Scenario Forced -- Control Std. Err. df $t$ ratio Adj. $P^{\dagger}$ Sig.
True $\times$ False 0.015 0.020 21428 0.756 1.000
True $\times$ Unsure 0.045 0.011 21428 3.992 $<$ 0.001 ***
True $\times$ True 0.091 0.023 21428 3.922 $<$ 0.001 ***
False $\times$ False 0.047 0.010 21428 4.943 $<$ 0.001 ***
False $\times$ Unsure 0.078 0.028 21428 2.739 0.0310 *
@@
$\dagger$ Bonferroni’s method comparing a family of 5 estimates
@@ 3.3 Opt in versus opt out
To provide support for our analysis related to the LLM-optional condition, we now incorporate an interaction between whether a participant in this condition chose to see LLM fact-checking information (opt in) or not (opt out) and the fact checking scenario.
Tables S9 and S10 present the results of fitting our data for the belief and share groups, respectively. To confirm the results presented in the main text, we utilize the models to perform the same comparisons of estimated marginal means. These post-hoc comparisons further support our findings, and are shown for the belief and share groups in Tables S11 and S12, respectively.
Table S9: Opt In versus Opt Out Coefficients (Belief Group; $F=286.42$, $R^{2}=0.23$, $P<0.001$)
@@
(Intercept) 0.628 0.061 10.251 $<0.001$**
Option(opt out) -0.243 0.039 -6.294 $<0.001$**
FC Scen.(False $\times$ unsure) 0.082 0.069 1.196 0.232
FC Scen.(True $\times$ false) 0.052 0.019 2.704 0.007 **
FC Scen.(True $\times$ true) 0.409 0.031 13.082 $<0.001$**
FC Scen.(True $\times$ unsure) 0.301 0.037 8.119 $<0.001$**
Age -0.004 0.001 -3.825 $<0.001$**
Education -0.008 0.009 -0.841 0.400
Option(opt out):FC Scen.(False $\times$ unsure) $<$ 0.000 0.072 0.002 0.999
Option(opt out):FC Scen.(True $\times$ false) 0.442 0.043 10.380 $<0.001$**
Option(opt out):FC Scen.(True $\times$ true) 0.216 0.046 4.673 $<0.001$**
Option(opt out):FC Scen.(True $\times$ unsure) 0.223 0.053 4.212 $<0.001$**
@@
Table S10: Opt In versus Opt Out Coefficients (Share Group; $F=217.95$, $R^{2}=0.19$, $P<0.001$)
@@
(Intercept) 0.805 0.067 12.066 $<0.001$**
Option(opt out) -0.305 0.039 -7.873 $<0.001$**
FC Scen.(False $\times$ unsure) 0.017 NaN
FC Scen.(True $\times$ false) -0.016 0.005 -2.905 0.004 **
FC Scen.(True $\times$ true) 0.104 0.021 5.022 $<0.001$**
FC Scen.(True $\times$ unsure) 0.072 0.018 4.128 $<0.001$**
Age -0.007 0.001 -5.258 $<0.001$**
Education 0.001 0.014 0.053 0.958
Option(opt out):FC Scen.(False $\times$ unsure) -0.007 0.013 -0.543 0.587
Option(opt out):FC Scen.(True $\times$ false) 0.065 0.021 3.104 0.002 **
Option(opt out):FC Scen.(True $\times$ true) -0.039 0.029 -1.348 0.178
Option(opt out):FC Scen.(True $\times$ unsure) 0.016 0.026 0.598 0.550
@@
Headline scenario Opt in -- Opt out Std. Error df $t$ ratio Adj. $P^{\dagger}$ Sig.
True $\times$ False -0.200 0.027 10428 -7.288 $<$ 0.001 ***
True $\times$ Unsure 0.020 0.015 10428 1.277 1.000
True $\times$ True 0.027 0.032 10428 0.837 1.000
False $\times$ False 0.243 0.013 10428 18.414 $<$ 0.001 ***
False $\times$ Unsure 0.243 0.039 10428 6.226 $<$ 0.001 ***
@@
$\dagger$ Bonferroni’s method comparing a family of 5 estimates
@@
Headline scenario Opt in -- Opt out Std. Error df $t$ ratio Adj. $P^{\dagger}$ Sig.
True $\times$ False 0.239 0.028 10508 8.501 $<$ 0.001 ***
True $\times$ Unsure 0.289 0.016 10508 18.347 $<$ 0.001 ***
True $\times$ True 0.344 0.032 10508 10.630 $<$ 0.001 ***
False $\times$ False 0.305 0.013 10508 22.876 $<$ 0.001 ***
False $\times$ Unsure 0.312 0.039 10508 7.917 $<$ 0.001 ***
@@
$\dagger$ Bonferroni’s method comparing a family of 5 estimates
@@ Appendix 4 Interaction analyses
In this section, we explore the potential moderation effects of two factors on our main results: attitude towards AI (ATAI) and headline congruence (see the Covariates section for details). We employ linear regression with robust standard errors clustered on participant and headline for each key finding discussed in the main text. Each analysis covered in the Regression analyses section is revisited to incorporate these variables and create three-way interactions. Covariates that were included in the earlier analyses (Age and Education level) are included again. The belief and sharing group data are modeled separately.
@@ 4.1 Attitude towards AI
We begin by examining whether LLM fact-checking information remains ineffective amongst individuals with varying levels of ATAI. Therefore, we test the three-way interaction between Condition, Veracity, and ATAI (Condition $\times$ Veracity $\times$ ATAI). The human fact checking group is excluded from this analysis, as there is no reason to believe that participants’ interactions with human-generated fact checks would vary based on their attitudes toward artificial intelligence. Figure S1 presents the relationship between participants’ ATAI and their belief in (panel a) and intent to share (panel b) true versus false headlines across all conditions. The results of our modeling analysis indicate that there is no significant three-way interaction between ATAI and either belief in (Table S13) or intent to share (Table S14) headlines for all conditions.
@@
Condition(Forced):Veracity(True) 0.027 0.085 0.315 0.753
@@
Condition(Optional):Veracity(True) 0.087 0.066 1.322 0.186
@@
Next, we examine whether the effects of fact-checking scenarios stay consistent among people with different ATAI (the three-way interaction Condition $\times$ FC Scenario $\times$ ATAI). Again, we focus on the forced and control conditions and exclude data for the optional participants when fitting each model. Figure S2 illustrates the relationship between belief in headlines and ATAI for the control and forced conditions in each fact-checking scenario. The same relationship is presented with respect to sharing intent in Figure S3. The result of fitting the belief and share group models are found in Tables S15 and S16, respectively. These models are then utilized for post-hoc comparisons similar to those presented within the main text for each group. However, to test for an ATAI interaction, this analysis compares the slopes of the Control and Forced groups predicted response line, given different values of ATAI. These results, shown in Tables S17 and S18 for the belief and share groups, respectively, validate our results by illustrating that participants did not respond differently depending on ATAI.
@@
Headline Scenario Forced -- Control Std. Err. df $t$ ratio Adj. $P^{\dagger}$ Sig.
False $\times$ false 0.010 0.008 19498 1.234 1.000
False $\times$ unsure 0.023 0.023 19498 1.000 1.000
True $\times$ false -0.038 0.016 19498 -2.313 0.104
True $\times$ unsure $<$ 0.001 0.009 19498 0.013 1.000
True $\times$ true 0.013 0.019 19498 0.658 1.000
@@
$\dagger$ Bonferroni’s method comparing a family of 5 estimates
@@
Headline Scenario Forced -- Control Std. Err. df $t$ ratio Adj. $P^{\dagger}$ Sig.
False $\times$ false -0.013 0.008 21418 -1.537 0.622
False $\times$ unsure 0.012 0.025 21418 0.495 1.000
True $\times$ false $<$ 0.001 0.018 21418 -0.017 1.000
True $\times$ unsure -0.014 0.010 21418 -1.449 0.736
True $\times$ true 0.015 0.021 21418 0.746 1.000
@@
$\dagger$ Bonferroni’s method comparing a family of 5 estimates
Next, we examine whether behavior in the optional condition depends on ATAI by introducing a three-way interaction term involving whether a participant chose to view LLM fact checks (opt in vs. opt out), fact checking scenario, and individual attitude towards AI (Opt-Condition $\times$ FC Scenario $\times$ ATAI). The results of fitting these models for the belief and share groups are presented in Tables S19 and S20, respectively. We conduct a post-hoc analysis that compares the slopes of the opt-in and opt-out conditions across varying levels of ATAI for the belief (Table S21) and sharing (Table S22) groups, respectively. Results of the post-hoc comparisons can be found in Tables S23 and S24 for the belief and sharing groups, respectively.
@@
[…]
Table S19: Opt In versus Opt Out Coefficients (ATAI interaction, Belief Group; $F=151.22.53$, $R^{2}=0.23$, $P<0.001$)
@@
(Intercept) 0.593 0.125 4.735 $<0.001$**
Option(opt out) -0.190 0.123 -1.550 0.121
FC Scen.(False $\times$ unsure) 0.081 NaN
FC Scen.(True $\times$ false) 0.142 0.053 2.647 0.008 **
FC Scen.(True $\times$ true) 0.425 0.130 3.271 0.001 ***
FC Scen.(True $\times$ unsure) 0.345 0.084 4.126 $<0.001$**
ATAI 0.008 0.023 0.347 0.729
Age -0.004 0.001 -3.907 $<0.001$**
Education -0.007 0.009 -0.714 0.475
Option(opt out):FC Scen.(False $\times$ unsure) -0.060 0.021 -2.906 0.004 **
Option(opt out):FC Scen.(True $\times$ false) 0.508 0.101 5.046 $<0.001$**
Option(opt out):FC Scen.(True $\times$ true) 0.182 0.221 0.822 0.411
Option(opt out):FC Scen.(True $\times$ unsure) 0.303 0.109 2.788 0.005 **
Option(opt out):ATAI -0.011 0.026 -0.424 0.671
FC Scen.(False $\times$ unsure):ATAI $<$ 0.001 NaN
FC Scen.(True $\times$ false):ATAI -0.019 0.007 -2.712 0.007 **
FC Scen.(True $\times$ true):ATAI -0.003 0.027 -0.122 0.903
FC Scen.(True $\times$ unsure):ATAI -0.009 0.016 -0.593 0.553
Option(opt out):FC Scen.(False $\times$ unsure):ATAI 0.015 NaN
Option(opt out):FC Scen.(True $\times$ false):ATAI -0.018 0.021 -0.836 0.403
Option(opt out):FC Scen.(True $\times$ true):ATAI 0.008 0.044 0.173 0.863
Option(opt out):FC Scen.(True $\times$ unsure):ATAI -0.020 0.022 -0.932 0.351
@@
(Intercept) 0.622 0.144 4.309 $<0.001$**
Option(opt out) -0.165 0.143 -1.154 0.248
FC Scen.(False $\times$ unsure) -0.080 NaN
FC Scen.(True $\times$ false) -0.017 NaN
FC Scen.(True $\times$ true) 0.082 0.082 0.991 0.322
FC Scen.(True $\times$ unsure) 0.045 0.065 0.686 0.493
ATAI 0.038 0.027 1.418 0.156
Age -0.007 0.001 -5.057 $<0.001$**
Education $<0.001$ 0.014 -0.021 0.983
Option(opt out):FC Scen.(False $\times$ unsure) 0.131 NaN
Option(opt out):FC Scen.(True $\times$ false) 0.067 0.047 1.436 0.151
Option(opt out):FC Scen.(True $\times$ true) 0.008 0.082 0.098 0.922
Option(opt out):FC Scen.(True $\times$ unsure) 0.080 0.093 0.859 0.391
Option(opt out):ATAI -0.029 0.030 -0.943 0.346
FC Scen.(False $\times$ unsure):ATAI 0.021 NaN
FC Scen.(True $\times$ false):ATAI $<0.001$ NaN
FC Scen.(True $\times$ true):ATAI 0.005 0.017 0.321 0.748
FC Scen.(True $\times$ unsure):ATAI 0.006 0.014 0.464 0.642
Option(opt out):FC Scen.(False $\times$ unsure):ATAI -0.031 NaN
Option(opt out):FC Scen.(True $\times$ false):ATAI -0.001 0.013 -0.053 0.958
Option(opt out):FC Scen.(True $\times$ true):ATAI -0.012 0.019 -0.618 0.537
Option(opt out):FC Scen.(True $\times$ unsure):ATAI -0.015 0.020 -0.772 0.440
@@
Option Headline Scenario $b$ Std. Err. df $t$-ratio $P$ Sig.
Opt in False $\times$ false 0.008 0.007 10418 1.200 0.230
Opt out False $\times$ false -0.003 0.007 10418 -0.437 0.662
Opt in False $\times$ unsure 0.008 0.019 10418 0.440 0.660
Opt out False $\times$ unsure 0.012 0.023 10418 0.512 0.609
Opt in True $\times$ false -0.011 0.013 10418 -0.817 0.414
Opt out True $\times$ false -0.040 0.016 10418 -2.464 0.014 **
Opt in True $\times$ true 0.005 0.016 10418 0.307 0.759
Opt out True $\times$ true 0.001 0.018 10418 0.068 0.946
Opt in True $\times$ unsure -0.001 0.007 10418 -0.166 0.868
Opt out True $\times$ unsure -0.033 0.009 10418 -3.552 $<$ 0.001 ***
@@
Option Headline Scenario $b$ Std. Err. df $t$-ratio $P$ Sig.
Opt in False $\times$ false 0.038 0.007 10498 5.269 $<$ .001 ***
Opt out False $\times$ false 0.009 0.008 10498 1.183 0.237
Opt in False $\times$ unsure 0.059 0.021 10498 2.789 0.005 **
Opt out False $\times$ unsure -4.71$\times$ 10-5 0.026 10498 -0.002 0.999
Opt in True $\times$ false 0.038 0.015 10498 2.665 0.008 **
Opt out True $\times$ false 0.009 0.019 10498 0.508 0.612
Opt in True $\times$ true 0.043 0.017 10498 2.560 0.011 *
Opt out True $\times$ true 0.003 0.021 10498 0.181 0.856
Opt in True $\times$ unsure 0.044 0.008 10498 5.497 $<$ .001 ***
Opt out True $\times$ unsure 0.001 0.010 10498 0.108 0.914